# Optimal Control of an SIR Model with Noncompliance as a Social Contagion

### v1

@@ 4.1. Scenario 1
The parameters for this scenario are listed in figure 1, and the result of the simulation in the absence of controls is included in figure 2. With no controls, the reproductive ratio for this scenario is $\mathscr{R}_{0}=\tfrac{b}{\delta}\cdot\frac{\beta}{\gamma+\delta}=\frac{0.4}{0.21}\approx 1.9.$ We note in particular that $\xi=0$, meaning that all newly introduced members of the population are compliant, and we are in the regime of theorem 2.3(i),(ii). In this case, there are two DFEs, that, when controls are constant, take the forms $(s_{1},s^{*}_{1})=(\tfrac{b}{\delta},0)$ and $(s_{2},s_{2}^{*})=(\tfrac{\nu+\delta}{\overline{\mu}-\mu},\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\overline{\mu}-\mu})$. If the controls are constant, the second is stable in the case that it is physically meaningful (i.e. $\tfrac{b}{\delta}<\tfrac{\nu+\delta}{\overline{\mu}-\mu}$); otherwise, the first is stable. Clearly to minimize (C1) and (C2), it behooves the policy-maker to drive the system toward $(s_{1},s_{1}^{*})=(\tfrac{b}{\delta},0)$. If control cost is no issue, the policy-maker can increase $\mu,\nu$ to ensure that $(s_{1},s_{1}^{*})$ is stable, and then decrease $\mathscr{R}_{0}(\tfrac{b}{\delta},0)=\tfrac{b}{\delta}\cdot\tfrac{\beta(1-\alpha)}{\gamma+\eta+\delta}$ using $\alpha,\mu$ so as to fall into the stability regime of theorem 2.3(i). However, it will likely be too expensive in terms of (C3) to actually fully do this. In figure 3, we see the optimal control plans for a “balanced” policymaker: one who would like to minimize both infections and noncompliance. The disease dynamics are plotted in the top panels figure 3, split into compliant and noncompliant compartments. The bottom left panel of figure 3 displays the optimal control plans. Finally, the bottom right displays the reproductive ratio $\mathscr{R}_{0}(s,s^{*})$ which, due to the controls, now changes with time. In particular, we plot the value of $\mathscr{R}_{0}(\tfrac{b}{\delta},0)$ when $\tfrac{b}{\delta}<\frac{\nu+\delta}{\overline{\mu}-\mu}$ to reflect the conditions of theorem 2.3(i) and $\mathscr{R}_{0}(\tfrac{\nu+\delta}{\overline{\mu}-\mu},\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\overline{\mu}-\mu})$ when $\tfrac{b}{\delta}\geq\frac{\nu+\delta}{\overline{\mu}-\mu}$ to reflect the conditions of theorem 2.3(ii). In either case, if the corresponding reproductive ratio was kept below 1, the corresponding DFE would be stable, but from the plot we see that it is not optimal to do so. Instead, controls $\mu$ and $\nu$ are used to ensure compliance in the early stages of the epidemic, so that $\alpha$ and $\eta$ can be effectively deployed to keep the reproductive ratio low, though a small outbreak is indeed allowed to occur. With these cost weights, the optimal controls reduce the total cost from $12.42$ in the uncontrolled case to $2.97$ in the optimally controlled case, representing a roughly $76\%$ relative reduction in cost due to controls.
@@
Figure 3. Simulation of Scenario 1 with a “balanced” policy-maker: cost weights $c_{1}=1,c_{2}=0.1$. The strategy is to quell noncompliance using $\mu,\nu$ so as to keep the infection rate low. Since the majority of the population is compliant, $\alpha$ and $\nu$ can be deployed to effectively lower the reproductive ratio. However, it is not lowered so much that $\mathscr{R}_{0}(\tfrac{b}{\delta},0)<1.$ A small outbreak of the disease is allowed to occur.
@@
Figure 4. Simulation of Scenario 1 with a public health oriented policy-maker: cost weights $c_{1}=5,c_{2}=0$. In this case, even though there is no explicit cost associated with noncompliance, because the policy-maker desires to keep infections low, the optimal strategy employs $\mu$ and $\nu$ even more strongly than in the balanced case. The disease controls $\alpha$ and $\eta$ are also increased so as to drive $\mathscr{R}_{0}(\tfrac{b}{\delta},0)<1$ and take advantage of the stability of the DFE $(\tfrac{b}{\delta},0)$ until the disease has been effectively eradicated.
@@
We contrast this with the optimal actions corresponding to cost weights $c_{1}=5,c_{2}=0$. This represents a public health oriented policy-maker whose primary concern is to stop the disease, and does not even assess cost of noncompliance. The results of the simulation using these weights are in figure 4. In this case, the cost weight on infections is high enough that it is indeed optimal to use controls $\mu,\nu$ to ensure that, in the initial stages of the epidemic, we remain in the stability regime for the disease-free equilibrium $(s_{1},s_{1}^{*})=(\tfrac{b}{\delta},0)$, and then use $\alpha,\eta$ to drive $\mathscr{R}_{0}(\tfrac{b}{\delta},0)<1$, in essense taking advantage of theorem 2.3(i). This is maintained until the disease has been effectively eradicated, at which time the policy-maker eases the controls (note that, in the context of this model, the policy-maker is aware of the horizon time $T=100$, so $\alpha$ is tapered off such that infections will not significantly increase within this final time). Figure 4 demonstrates the need for synergy between the disease controls $\alpha,\eta$ and noncompliance controls $\mu,\nu$. With these cost weights, there is no explicit cost associated with increased noncompliance. However, because increased noncompliance accelerates disease spread, the policy-maker must still employ $\mu,\nu$ to keep noncompliance low so as to accomplish their goal of disease elimination. With these cost weights, the total cost is reduced from $22.60$ in the uncontrolled case to $6.90$ in the optimally controlled case, representing a roughly $70\%$ relative reduction in cost.
@@
To further explore the synergy between the controls, we include one final figure using the scenario 1 parameters, and the health oriented policy-maker with cost weights $c_{1}=5,c_{2}=0$. In this case, we demonstrate the effects of “turning off” certain controls. The results are in figure 5, where we display the effects on the optimal control maps of forcing $\eta=0$ (top), or forcing $\mu=\nu=0$ (bottom). Recall, $\eta$ represents treatment efforts that hasten recovery among the compliant infected population, so one could think of $\eta=0$ as a scenario near the outset of an epidemic where treatment is unavailable or ineffective. In this case, the policy-maker’s optimal strategy is to more strongly enforce NPIs (increase $\alpha$), while also using $\mu,\nu$ to ensure compliance. While they are not displayed, the dynamics in this case look, in essence, the same as those in the top panel of figure 4 in that the disease is effectively eradicated. However, without access to the control $\eta$, the total cost can only be reduced from $22.60$ in the uncontrolled case to $15.43$ in the optimally controlled case, representing only a roughly $32\%$ relative cost reduction (compared to a $70\%$ relative cost reduction when $\eta$ is available). Thus, in some sense, controls efforts are only half as effective without $\eta$.
@@
Figure 5. Simulation of Scenario 1 with a health oriented policy-maker: cost weights $c_{1}=5,c_{2}=0$, with either treatment $\eta$ unavailable (top), or control of noncompliance $\mu,\nu$ unavailable (bottom). Note the $y$-axis change in the top left.
@@
In a similar vein, $\mu=\nu=0$ represents a scenario wherein it is impossible to encourage compliance or slow the spread of noncompliance. In this case, the dynamics look in essence like those in figure 2 where no controls are used. One sees in the bottom panel of figure 5 that without the ability to curb noncompliance, it is still beneficial to slightly decrease the reproductive ratio among the compliant population using $\alpha$ and $\eta$, but much less so since the majority of the population will become noncompliant, hampering disease prevention efforts. With $\mu=\nu=0$, one cannot land in the stability regime of theorem 2.3(i), nor accomplish $\mathscr{R}_{0}(\frac{\nu+\delta}{\overline{\mu}-\mu},\frac{b}{\delta}-\frac{\nu+\delta}{\overline{\mu}-\mu})<1$, so the disease is simply allowed to run its course. The use of $\alpha$ and $\eta$ still accomplish a reduction of total cost from $22.60$ in the uncontrolled case to $19.69$ in the optimally controlled case, representing a roughly $13\%$ relating reduction in total cost (compared to the $70\%$ cost reduction when all controls are available). Meaning that, with $\mu=\nu=0$, controls efforts are only about one-fifth as effective.
@@ 4.2. Scenario 2
In this scenario, when compared with scenario 1, we increase the disease infection rate $\beta$ and the spread rate $\overline{\mu}$ of noncompliance, and we also assume that $30\%$ of the newly introduced population members are noncompliant ($\xi=0.3$), while all other parameters are unchanged. For this scenario, we set $c_{2}=0.01$, indicating a weak preference of the policy-maker to quell noncompliance for its own sake, though quelling noncompliance for the sake of slowing disease spread is still desirable. In figure 6, we display the results of the simulation with $c_{1}$ successively taking values $1/3,1,3,6,9$. In this case, noncompliance is much more difficult to control since 30% of the newly introduced population is noncompliant, and noncompliance spreads fairly quickly. When noncompliance is difficult to control, the control variables $\alpha$ and $\eta$ have a diminished effect. Accordingly, in this parameter regime, even a strongly public-health oriented policy-maker will not fully suppress disease spread. Note, for example, that the dynamics are qualitatively similar in the cases of $c_{1}=1$ and $c_{1}=6$. It isn’t until $c_{1}=9$ that the policy-maker significantly increases the use of the control $\alpha$ so as to more effectively stunt disease spread. This indicates that in scenarios where disease spread is fast enough, and noncompliance is rampant enough, a policy-maker who is even slightly economically minded (i.e. concerned with keeping control costs down) will let the disease run its course.
Figure 6. Simulation of Scenario 2, with increased spread of the disease an noncompliance. Noncompliance cost weight is $c_{2}=0.01$. From top to bottom, the cost weight on infections is $c_{1}=1/3,1,3,6,9$. Because disease spread is so fast, suppression efforts are costly and it is not optimal to fully suppress the disease until $c_{1}$ is near 99. In a case like this, an economically oriented policy-maker would allow the disease to spread.
@@ 4.3. Scenario 3
Finally, in this scenario, we revert to the parameters of the baseline simulation in figure 3, except that we increase the spread rate for noncompliance to $\overline{\mu}=0.5$ and observe the effects of changing $\xi$, the proportion of the newly introduced population that is noncompliant. In particular, we have the “balanced” cost weights $c_{1}=1,c_{2}=0.1$. The full parameter list for scenario 3 is in figure 1. In some sense, $\xi$ could be interpreted as the baseline proclivity toward noncompliance within the population. A key question is then the level at which this proclivity is high enough that it is no longer optimal to attempt to suppress noncompliance. We include four plots in figure 7 where $\xi$ takes successive values $0,1/4,1/2,1.$ When $\xi=0$, even with the very high rate of noncompliance spread, it is optimal to fully quell noncompliance in the early stages of the epidemic. This is seen, for example, in the top panels of figure 7, where $\mu$ still takes its maximum value for a short time at the beginning, and noncompliance stays throughout most of the time interval (we note again that it only grows at the end because the policy-maker is aware of the horizon time, and can save some cost by releasing some controls when there is no longer time for an outbreak to occur). This is contrasted with the ensuing plots when $\xi=1/4,1/2,1$. In those cases, $\mu$ never takes it maximum value, and when $\xi=1/2,1$, $\mu$ remains below half its maximum value. The results for the cases $\xi=1/2$ and $\xi=1$ are very similar, seeming to indicate that for values $\xi\geq 1/2$, controls are simply ineffective. We note also that $\alpha$ and $\nu$ decrease with increasing $\xi$, which is a result of the fact that disease control is less effective when control of noncompliance cannot be achieved.
Figure 7. Simulation of Scenario 3, with increased spread rate of noncompliance, balanced cost weights, and variable proportion of the newly introduced population which is noncompliant $(\xi)$. Even with the increased spread of noncompliance, as long as $\xi=0$, the optimal strategy still employs the maximum available $\mu$ for a small time to suppress spread of noncompliance. As $\xi$ increases, slowing spread of noncompliance becomes futile, so that even when $\xi=0.25$ the policy-maker no longer uses the maximum value of $\mu$, and when $\xi=1/2$, $\mu$ is never more than half its maximum value and essentially the entire epidemic occurs on the noncompliant side. There is very little change between the simulations using $\xi=1/2$ and $\xi=1$, indicating that by the time $\xi\geq 1/2$, a threshold has been crossed and controls are ineffective.

### v2

@@ 5.1. Scenario 1
The parameters for this scenario are listed in figure 1, and the result of the simulation in the absence of controls is included in figure 2. With no controls, the reproductive ratio for this scenario is $\mathscr{R}_{0}=\tfrac{b}{\delta}\cdot\frac{\beta}{\gamma+\delta}=\frac{0.4}{0.21}\approx 1.9.$ We note in particular that $\xi=0$, meaning that all newly introduced members of the population are compliant, and we are in the regime of theorem 2.3(i),(ii). In this case, there are two DFEs, that, when controls are constant, take the forms $(s_{1},s^{*}_{1})=(\tfrac{b}{\delta},0)$ and $(s_{2},s_{2}^{*})=(\tfrac{\nu+\delta}{\mu^{*}-\mu},\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\mu^{*}-\mu})$. If the controls are constant, the second is stable in the case that it is physically meaningful (i.e. $\tfrac{b}{\delta}<\tfrac{\nu+\delta}{\mu^{*}-\mu}$); otherwise, the first is stable. Clearly to minimize (C1) and (C2), it behooves the policy-maker to drive the system toward $(s_{1},s_{1}^{*})=(\tfrac{b}{\delta},0)$. If control cost is no issue, the policy-maker can increase $\mu,\nu$ to ensure that $(s_{1},s_{1}^{*})$ is stable, and then decrease $\mathscr{R}_{0}(\tfrac{b}{\delta},0)=\tfrac{b}{\delta}\cdot\tfrac{\beta(1-\alpha)}{\gamma+\eta+\delta}$ using $\alpha,\eta$ so as to fall into the stability regime of theorem 2.3(i). However, it will likely be too expensive in terms of (C3) to actually fully do this. In figure 3, we see the optimal control plans for a “balanced” policymaker: one who would like to minimize both infections and noncompliance. The disease dynamics are plotted in the top panels figure 3, split into compliant and noncompliant compartments. The bottom left panel of figure 3 displays the optimal control plans. Finally, the bottom right displays the reproductive ratio $\mathscr{R}_{0}(s,s^{*})$ which, due to the controls, now changes with time. In particular, we plot the value of $\mathscr{R}_{0}(\tfrac{b}{\delta},0)$ when $\tfrac{b}{\delta}<\frac{\nu+\delta}{\mu^{*}-\mu}$ to reflect the conditions of theorem 2.3(i) and $\mathscr{R}_{0}(\tfrac{\nu+\delta}{\mu^{*}-\mu},\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\mu^{*}-\mu})$ when $\tfrac{b}{\delta}\geq\frac{\nu+\delta}{\mu^{*}-\mu}$ to reflect the conditions of theorem 2.3(ii). In either case, if the corresponding reproductive ratio was kept below 1, the corresponding DFE would be stable, but from the plot we see that it is not optimal to do so. Instead, controls $\mu$ and $\nu$ are used to ensure compliance in the early stages of the epidemic, so that $\alpha$ and $\eta$ can be effectively deployed to keep the reproductive ratio low, though a small outbreak is indeed allowed to occur. With these cost weights, the optimal controls reduce the total cost from $8.47$ in the uncontrolled case to $4.79$ in the optimally controlled case, representing a roughly $43.5\%$ relative reduction in cost due to controls.
@@
Figure 3. Simulation of Scenario 1 with a “balanced” policy-maker: cost weights $c_{1}=1,c_{2}=0.05$, $d_{1}=0.1,d_{2}=0.1$. The strategy is to quell noncompliance using $\mu,\nu$ so as to keep the infection rate low. Since the majority of the population is compliant, $\alpha$ and $\nu$ can be deployed to effectively lower the reproductive ratio. However, it is not lowered so much that $\mathscr{R}_{0}(\tfrac{b}{\delta},0)<1.$ A small outbreak of the disease is allowed to occur.
@@
Figure 4. Simulation of Scenario 1 with a public health oriented policy-maker: cost weights $c_{1}=5,c_{2}=0,d_{1}=0.1,d_{2}=0.1$. In this case, even though there is no explicit cost associated with noncompliance, because the policy-maker desires to keep infections low, the optimal strategy employs $\mu$ and $\nu$ even more strongly than in the balanced case. The disease controls $\alpha$ and $\eta$ are also increased so as to drive $\mathscr{R}_{0}(\tfrac{b}{\delta},0)<1$ and take advantage of the stability of the DFE $(\tfrac{b}{\delta},0)$ until the disease has been effectively eradicated.
@@
We contrast this with the optimal actions corresponding to cost weights $c_{1}=5,c_{2}=0$. This represents a public health oriented policy-maker whose primary concern is to stop the disease, and does not even assess cost based on noncompliance. The results of the simulation using these weights are in figure 4. In this case, the cost weight on infections is high enough that it is indeed optimal to use controls $\mu,\nu$ to ensure that, in the initial stages of the epidemic, we remain in the stability regime for the disease-free equilibrium $(s_{1},s_{1}^{*})=(\tfrac{b}{\delta},0)$, and then use $\alpha,\eta$ to drive $\mathscr{R}_{0}(\tfrac{b}{\delta},0)<1$, in essence taking advantage of theorem 2.3(i). This is maintained until the disease has been effectively eradicated, at which time the policy-maker eases the controls (note that, in the context of this model, the policy-maker is aware of the horizon time $T=100$, so $\alpha$ is tapered off such that infections will not significantly increase within this final time). Figure 4 demonstrates the need for synergy between the disease controls $\alpha,\eta$ and noncompliance controls $\mu,\nu$. With these cost weights, there is no explicit cost associated with increased noncompliance. However, because increased noncompliance accelerates disease spread, the policy-maker must still employ $\mu,\nu$ to keep noncompliance low so as to accomplish their goal of disease elimination. With these cost weights, the total cost is reduced from $22.60$ in the uncontrolled case to $8.62$ in the optimally controlled case, representing a roughly $61.9\%$ relative reduction in cost.
@@
To further explore the synergy between the controls, we include one final figure using the scenario 1 parameters, and the health oriented policy-maker with cost weights $c_{1}=5,c_{2}=0$. In this case, we demonstrate the effects of “turning off” certain controls. The results are in figure 6, where we display the effects on the optimal control maps of forcing $\eta=0$ (top), or forcing $\mu=\nu=0$ (bottom). Recall, $\eta$ represents treatment efforts that hasten recovery among the compliant infected population, so one could think of $\eta=0$ as a scenario near the outset of an epidemic where treatment is unavailable or ineffective. In this case, the policy-maker’s optimal strategy is to more strongly enforce NPIs (increase $\alpha$), while also using $\mu,\nu$ to ensure compliance. While they are not displayed, the dynamics in this case look, in essence, the same as those in the top panel of figure 4 in that the disease is effectively eradicated. However, without access to the control $\eta$, the total cost can only be reduced from $22.60$ in the uncontrolled case to $13.16$ in the optimally controlled case, representing only a roughly $41.8\%$ relative cost reduction (compared to a $61.9\%$ relative cost reduction when $\eta$ is available). Thus, in some sense, controls efforts are only two-thirds as effective without $\eta$.
@@
Figure 5. Simulation of Scenario 1 with a health oriented policy-maker: cost weights $c_{1}=5,c_{2}=0,d_{1}=0.1,d_{2}=0.1$, with either treatment $\eta$ unavailable (top), or control of noncompliance $\mu,\nu$ unavailable (bottom). Note the $y$-axis change in the top left.
@@
In a similar vein, $\mu=\nu=0$ represents a scenario wherein it is impossible to encourage compliance or slow the spread of noncompliance. In this case, the dynamics look in essence like those in figure 2 where no controls are used. One sees in the bottom panel of figure 6 that without the ability to curb noncompliance, it is still beneficial to slightly decrease the reproductive ratio among the compliant population using $\alpha$ and $\eta$, but much less so since the majority of the population will become noncompliant, hampering disease prevention efforts. With $\mu=\nu=0$, one cannot land in the stability regime of theorem 2.3(i), nor accomplish $\mathscr{R}_{0}(\frac{\nu+\delta}{\mu^{*}-\mu},\frac{b}{\delta}-\frac{\nu+\delta}{\mu^{*}-\mu})<1$, so the disease is simply allowed to run its course. The use of $\alpha$ and $\eta$ still accomplish a reduction of total cost from $22.60$ in the uncontrolled case to $20.31$ in the optimally controlled case, representing a roughly $10.1\%$ relative reduction in total cost (compared to the $61.9\%$ cost reduction when all controls are available). Meaning that, with $\mu=\nu=0$, controls efforts are only about one-sixth as effective.
Figure 6. Simulation of Scenario 1 with a health oriented policy-maker: cost weights $c_{1}=5,c_{2}=0,d_{1}=0.1,d_{2}=0.1$, with either treatment $\eta$ unavailable (top), or control of noncompliance $\mu,\nu$ unavailable (bottom).
Finally, for our last simulation in scenario 1, we demonstrate that our numerical can handle the case of $L^{1}$ control costs by setting $d_{2}=0$, eliminating the $L^{2}$ costs from (5.1). In this case, the Hamiltonian (3.9) is affine in $\eta,\nu\mu$, while still being quadratic in $\alpha$, since the dynamics are quadratic in $\alpha$. Because an affine function on an interval maximized at one of the endpoints, in this case, $\eta,\mu,\nu$ should only ever take the value 0 or their maximum value; these are akin to bang-bang controls (though they could perhaps more accurately be called “bang-no bang”). The control $\alpha$ does not have this structure and can still vary smoothly. Indeed, this is precisely what we see in figure 7. This demonstrates that our methods are well-equipped to handle $L^{1}$ control problems. It also shows why allowing for controls in $BV([0,T];U_{\text{ad}})$ is far preferable to higher regularity spaces: controls which are discontinuous but have bounded variation readily appear in applications.
Figure 7. Simulation of Scenario 1 with a balanced policy-maker with no $L^{2}$ control costs: cost weights $c_{1}=1,c_{2}=0,d_{1}=0.1,d_{2}=0$. Theoretically, the controls $\eta,\mu,\nu$ should only take their maximum value or zero, and indeed this is borne out in the simulation. This demonstrates that our methods are well-equipped to handle $L^{1}$ control problems.
@@ 5.2. Scenario 2
In this scenario, when compared with scenario 1, we increase the disease infection rate $\beta$ and the spread rate $\mu^{*}$ of noncompliance, and we also assume that $30\%$ of the newly introduced population members are noncompliant ($\xi=0.3$), while all other parameters are unchanged. For this scenario, we set $c_{2}=0.01$, indicating a weak preference of the policy-maker to quell noncompliance for its own sake, though quelling noncompliance for the sake of slowing disease spread is still desirable. In figure 8, we display the results of the simulation with $c_{1}$ successively taking values $1/3,1,3,6,9$. In this case, noncompliance is much more difficult to control since 30% of the newly introduced population is noncompliant, and noncompliance spreads fairly quickly. When noncompliance is difficult to control, the control variables $\alpha$ and $\eta$ have a diminished effect. Accordingly, in this parameter regime, only a very strongly public-health oriented policy-maker will fully suppress disease spread. Note, for example, that the dynamics are qualitatively similar in the cases of $c_{1}=1$ and $c_{1}=3$. It isn’t until $c_{1}=6$ that the policy-maker significantly increases the use of the control $\alpha$ so as to more effectively stunt disease spread. This indicates that in scenarios where disease spread is fast enough, and noncompliance is rampant enough, a policy-maker who is even slightly economically minded (i.e. concerned with keeping control costs down) will let the disease run its course.
Figure 8. Simulation of Scenario 2, with increased spread of the disease an noncompliance. Noncompliance cost weight is $c_{2}=0.01$. From top to bottom, the cost weight on infections is $c_{1}=1/3,1,3,6,9$. Because disease spread is so fast, suppression efforts are costly and it is not optimal to fully suppress the disease until $c_{1}\geq 6$. In a case like this, an economically oriented policy-maker would allow the disease to spread.
@@ 5.3. Scenario 3
Finally, in this scenario, we revert to the parameters of the baseline simulation in figure 3, except that we increase the spread rate for noncompliance to $\mu^{*}=0.3$ and observe the effects of changing $\xi$, the proportion of the newly introduced population that is noncompliant. In particular, we have the “balanced” cost weights $c_{1}=1,c_{2}=0.05$. The full parameter list for scenario 3 is in figure 1. In some sense, $\xi$ could be interpreted as the baseline proclivity toward noncompliance within the population. A key question is then the level at which this proclivity is high enough that it is no longer optimal to attempt to suppress noncompliance. We include four plots in figure 9 where $\xi$ takes successive values $0,1/4,1/2,1.$ When $\xi=0$, even with the very high rate of noncompliance spread, it is optimal to significantly quell noncompliance in the early stages of the epidemic. This is seen, for example, in the top panels of figure 9, where $\mu$ still takes its maximum value for at the beginning, and noncompliance stays relatively low throughout most of the time interval (we note again that it only grows at the end because the policy-maker is aware of the horizon time, and can save some cost by releasing some controls when there is no longer time for an outbreak to occur). This behavior persists for $\xi=1/4$, but is contrasted with the ensuing plots when $\xi=1/2,1$. In those cases, $\mu$ is almost identically zero. The results int the cases $\xi=1/2$ and $\xi=1$ are very similar, seeming to indicate that for values $\xi\geq 1/2$, controls are simply ineffective. We note also that $\alpha$ and $\nu$ decrease with increasing $\xi$, which is a result of the fact that disease control is less effective when control of noncompliance cannot be achieved.
Figure 9. Simulation of Scenario 3, with increased spread rate of noncompliance, balanced cost weights, and variable proportion of the newly introduced population which is noncompliant $(\xi)$. Even with the increased spread of noncompliance, as long as $\xi=0,1/4$, the optimal strategy still employs the maximum available $\mu$ for some time to suppress spread of noncompliance becomes futile so that when $\xi\geq 1/2$, $\mu$ is essentially zero. There is very little change between the simulations using $\xi=1/2$ and $\xi=1$, indicating that by the time $\xi\geq 1/2$, a threshold has been crossed and controls are ineffective.